OpenStudy (summersnow8):

What is the pH of a solution that is prepared by mixing 100 mL of 0.50 M sodium acetate and 115 mL of 0.100 M acetic acid? @photon336

OpenStudy (daniellelovee):

@agent0smith (one of the best for science)

OpenStudy (jiteshmeghwal9):

\(NaOH+CH_3COOH \rightarrow \text{Solution}\) pH of a solution = \(-log[H^+]\)

OpenStudy (jiteshmeghwal9):

\(CH_3COOH \rightarrow CH_3COOH^- + H^+\) Initial concentration= 0.100M 0 0 Final concentration= 0.100(1-\(\alpha\) ) \(0.100\alpha\) \(0.100 \alpha\)

OpenStudy (jiteshmeghwal9):

\[K_a= \frac{(0.110 \alpha)(0.110 \alpha)}{0.110(1- \alpha)}\]where \(k_a\) is the dissociation constant for acetic acid & is equal to \(1.75 \times 10^{-5}\).

OpenStudy (jiteshmeghwal9):

from here u may find \([H^+]=-0.110 \alpha\) & then u may find pH of the solution

OpenStudy (summersnow8):

@jiteshmeghwal9 what does alpha refer to?

OpenStudy (kevin):

n acid = 115 x 0.100 = 11.5 mmol n base = 100 x 0.50 =50 mmol H+ = Ka *acid/ base H+ = 1.75 × 10 −5 * 11.5/50 H+ = 4 x 10^-6 pH = - log [H+] pH = - log 4 x 10^-6 pH = 5.40

OpenStudy (kevin):

@Summersnow8 What do u think?

OpenStudy (summersnow8):

where did you get the ka?

OpenStudy (kevin):

I'm searching...

OpenStudy (kevin):

We don't have [H+] So I think we can't find Ka

OpenStudy (kevin):

Should we find it using formula?

OpenStudy (summersnow8):

I am not sure. where did you get 1.75 × 10 −5

OpenStudy (kevin):

searching on google

OpenStudy (kevin):

It said acetic has 1.75 x 10^-5

OpenStudy (summersnow8):

hmmmmm, well that won't work. I need to be able to solve it without the computer

OpenStudy (kevin):

What formula we can use to find Ka?

OpenStudy (summersnow8):

Ka = ( [H+] [A¯] ) / [HA]

OpenStudy (kevin):

What is A- and HA ?

OpenStudy (kevin):

Sorry I'm asking since I really don't know what is it. In my country, maybe it written in other word.

OpenStudy (summersnow8):

A would be the base, while HA would be an acid

OpenStudy (kevin):

That's what we're looking for, the [H+] I will ask someone about this question, maybe he can help.

OpenStudy (summersnow8):

okay